# Computing the $q$-Multiplicity of the Positive Roots of $\mathfrak{sl}_{r+1}(\mathbb{C})$ and Products of Fibonacci Numbers

### v1

@@ Abstract
Using Kostant’s weight multiplicity formula, we describe and enumerate the terms contributing a nonzero value to the multiplicity of a positive root $\mu$ in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$, which we denote $L(\tilde{\alpha})$, where $\tilde{\alpha}$ is the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. We prove that the number of terms contributing a nonzero value in the multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in $L(\tilde{\alpha})$ is given by the product $F_{i}\cdot F_{r-j+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. Using this result, we show that the $q$-multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in the representation $L(\tilde{\alpha})$ is precisely $q^{r-h(\mu)}$, where $h(\mu)=j-i+1$ is the height of the positive root $\mu$. Setting $q=1$ recovers the known result that the multiplicity of a positive root in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$ is one.
@@ 1 Introduction
Let $G$ be a simple linear algebraic group over $\mathbb{C}$, $T$ be a maximal algebraic torus in $G$ of dimension $r$, and $B$ be a choice of a Borel subgroup where $T\subseteq B\subseteq G$. Let $\mathfrak{g}$, $\mathfrak{h}$, and $\mathfrak{b}$ be the Lie algebras of $G$, $T$, and $B$, respectively. Let the set of roots corresponding to $(\mathfrak{g},\mathfrak{h})$ be denoted as $\Phi$ and let the set of positive roots with respect to $\mathfrak{b}$ be denoted as $\Phi^{+}$. Note that $\Phi^{+}\subseteq\Phi$. Let the simple roots be denoted as $\Delta\subseteq\Phi^{+}$. We let the Weyl group corresponding to $G$ and $T$ be denoted as $W=Norm_{G}(T)/T$. Recall that the Weyl group is generated by reflections about hyperplanes lying perpendicular to the simple roots of the Lie algebra $\mathfrak{g}$. For any $\sigma\in W$, the length of $\sigma$, denoted $\ell(\sigma)$, is equal to the minimal nonnegative integer $k$ such that $\sigma$ is a product of $k$ simple reflections.
It is well-known that a finite-dimensional irreducible representation $V$ of a simple Lie algebra $\mathfrak{g}$ can be studied by decomposing
$V=\oplus V_{\alpha}$ (1)
where the direct sum is indexed by a finite set of weights, which are linear functionals from the Cartan subalgebra to $\mathbb{C}$. When given a weight $\mu$, the corresponding subspace $V_{\mu}$ is called a weight space and the dimension of $V_{\mu}$ is called the multiplicity of $\mu$. To study representations of $\mathfrak{g}$, it is sufficient to determine the multiplicity of the weights appearing in equation (1).
To compute the multiplicity of a weight $\mu$ in the irreducible representation with highest weight $\lambda$, we can use Kostant’s weight multiplicity formula [10]:
@@
where $\rho=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha$ and $\wp(\xi)$ is Kostant’s partition function. Recall that Kostant’s partition function $\wp(\xi)$ counts the number of ways the weight $\xi$ can be written as a nonnegative integral sum of positive roots.
When using equation (2), every element in the Weyl group contributes a term to the multiplicity. This forms a challenge since the order of the Weyl group increases factorially as the rank of the Lie algebra increases and many of the elements contribute trivially (a zero term) to the computation. In order to formulate a more efficient way to know what elements of the Weyl group contribute to Kostant’s weight multiplicity formula, we recall the following definition [6, Definition 1.1].
Definition 1.1.
For $\lambda,\mu$ integral weights of $\mathfrak{g}$, define the Weyl alternation set to be
@@
The Weyl alternation set tells us the elements of the Weyl group for which $\sigma(\lambda+\rho)-\mu-\rho$ can be written as a nonnegative integral sum of positive roots.
The importance of the Weyl alternation set is that it reduces the multiplicity computation to its support. On this smaller contributing set, it is often the case that we can give formulas for the $q$-analog of Kostant’s partition function. We recall the definition next.
Definition 1.2.
The $q$-analog of Kostant’s partition function is the polynomial-valued function defined by
$\wp_{q}(\xi)=c_{0}+c_{1}q+c_{2}q^{2}+\cdots+c_{k}q^{k},$
@@
where $c_{i}$ equals the number of ways to write $\xi$ as a sum of exactly $i$ positive roots.
@@
The work on enumerating and characterizing the Weyl alternation sets for certain pairs of weights, $(\lambda,\mu)$, has gathered some attention in the literature. This includes the following. If $\lambda$ and $\mu$ are integral weights of the Lie algebras $\mathfrak{sl}_{3}(\mathbb{C})$ and $\mathfrak{sl}_{4}(\mathbb{C})$, then the sets $\mathcal{A}(\lambda,\mu)$ are known in general [3, 11]. Harris considered $\tilde{\alpha}$, the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$ with $r\geq 1$, and $\mu=0$ and established the cardinality of the Weyl alternation set, $|\mathcal{A}(\tilde{\alpha},0)|$, is a Fibonacci number [6, Theorem 1.2]. Harris also stated that if $\mu$ is a dominant weight, then the Weyl alternation set is either $\{1\}$ when $\mu=\tilde{\alpha}$ or empty otherwise. Chang, Harris, and Insko extended these results and showed that when $\mu=0$ and $\lambda$ is the sum of all the simple roots of the simple Lie algebra $\mathfrak{so}_{2r+1}(\mathbb{C})$, the cardinality of the Weyl alternation set $\mathcal{A}(\lambda,0)$ is given by a Fibonacci number [2, Corollary 2.1 and 3.1]. While in the Lie algebras $\mathfrak{sp}_{2r}(\mathbb{C})$ and $\mathfrak{so}_{2r}(\mathbb{C})$, the cardinality of the Weyl alternation sets $\mathcal{A}(\lambda,0)$ are given by a multiple of a Lucas number [2, Corollary 4.1 and 5.1]. Harris, Insko, and Williams consider $\tilde{\alpha}$, the highest root of the classical Lie algebras, and $\mu=0$ and show that $|\mathcal{A}(\tilde{\alpha},0)|$ satisfies homogeneous linear recurrence relations [8]. In the type $B$ case, if $\mu=\varpi_{1}$ denotes the first fundamental root, then $|\mathcal{A}(\lambda,\varpi_{1})|$ is a Fibonacci number [8, Corollary 8.4]. Harris et al. also consider the case where $\mu$ is a dominant positive root and give the multiplicity of these weights in the adjoint representation of the classical Lie algebras. Berenshteĭn and Zelevinskiĭ [1, Theorem 1.3] provided a comprehensive list of pairs of weights $(\lambda,\mu)$ of the classical Lie algebras and the exceptional Lie algebra $G_{2}$ for which $m(\lambda,\mu)=1$. Harris, Rahmoeller, Schneider, and Simpson consider certain such pairs, determine $\mathcal{A}(\lambda,\mu)$, and show that the $q$-multiplicity is a power of $q$ [9]. For all pairs of weights that Berenshteĭn and Zelevinskiĭ provided, Harris et. al conjecture that $m_{q}(\lambda,\mu)$ is a power of $q$ [9, Conjecture 7].
@@
As noted above, past work fixed $\lambda$ to be the highest root or the sum of the simple roots of a classical Lie algebra, while $\mu$ was either 00 or a dominant positive root. In this work, we focus our study on $\mathfrak{sl}_{r+1}(\mathbb{C})$ and let $\Delta=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ be the set of simple roots and $\Phi^{+}=\Delta\cup\{\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}:1\leq i<j\leq r\}$ be the set of positive roots. Throughout all of our results, we fix $r\geq 1$, let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$, and consider $\mu\in\Phi^{+}$. In this way, $\mu$ may be a nondominant weight. Recall, that it is known that whenever $\mu$ is a root, $m(\tilde{\alpha},\mu)=1$. Hence, when $\mu$ is dominant, the pair $(\lambda,\mu)$ is given by Berenshteĭn and Zelevinskiĭ [1]. Moreover, we recall that the multiplicity is invariant under the action of the Weyl group. Namely, $m(\tilde{\alpha},\mu)=m(\tilde{\alpha},\sigma(\mu))$ for all $\sigma\in W$ [4, Proposition 3.1.20]. However, when $\mu\in\Phi^{+}$ is nondominant, a characterization and enumeration of the elements in $\mathcal{A}(\tilde{\alpha},\mu)$ as well as the value of $m_{q}(\tilde{\alpha},\mu)$ were unknown. Our work contributes this to the literature by proving the following results.
@@
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
@@
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
$|\mathcal{A}(\tilde{\alpha},\mu)|=F_{i}\cdot F_{r-j+1},$
[…]
Given the characterization and enumeration of the terms contributing to the $q$-multiplicity formula, we are now able to provide the following result.
@@
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
@@
Observe that setting $q=1$ in Theorem 4.1 recovers the known result that the multiplicity of a positive root in the adjoint representation is equal to one [4, Theorem 2.4.1].
This work is organized as follows: In Section 2, we present the proof to Theorem 2.1 showing that the cardinality of the Weyl alternation is a product of Fibonacci numbers. In Section 3, we establish results related to the $q$-multiplicity. In Section 4, we present the proof to Theorem 4.1 showing that the $q$-multiplicity of a positive root simplifies down to a power of $q$.
Acknowledgements. The author thanks her adviser Pamela E. Harris for introduction to the problem and feedback on an earlier draft. She also thanks Jeb Willenbring for insightful comments throughout the progression of this manuscript. The author gratefully acknowledge funding from the UW-Milwaukee Department of Mathematical Sciences in support of this research.
@@ 2 Weyl Alternation Sets
In [2], they consider the case where $\lambda$ is the highest root and $\mu=0$. We will still let $\lambda$ be the highest root and denote it as $\tilde{\alpha}$. However, we let $\mu\in\Phi^{+}$. The following result characterizes the elements in the set $\mathcal{A}(\tilde{\alpha},\mu)$.
@@
$\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
@@
($\Rightarrow$) We proceed by contradiction and assume $\sigma\neq 1$ or $\sigma\neq s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$. Then $\sigma$ must be either $s_{1}$, $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or must contain some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
@@
since $-\alpha_{1}+\alpha_{2}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r}$ cannot be written as a sum of positive roots since the coefficient for $\alpha_{1}$ is negative. Thus, we get a contradiction, and $s_{1}\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. By [8, Proposition 3.4], if $\sigma\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$, then neither is any $\sigma^{\prime}\in W$ that contains $\sigma$ in its reduced word expression. Thus, $\sigma\in W$ containing $s_{1}$ in its reduced word expression cannot be in $\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. This reaches a contradiction to the assumption that $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$.
A similar argument is made for when $\sigma$ is either $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or must contain some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
@@
Notice that $\{i,i+1,\ldots,j\}\cap\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}=\emptyset$. Now if $x\in\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (4), the coefficient of $\alpha_{x}$ is exactly 0. If $x\notin\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (4), the coefficient of $\alpha_{x}$ is exactly 1. In either case, $\tilde{\alpha}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})-\sum_{h=1}^{t}\alpha_{\ell_{h}}$, can be written as a sum of positive roots. Therefore, $s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. ∎
@@
We will utilize the following enumerative result in later results.
@@
[…]

### v2

@@ Abstract
Using Kostant’s weight multiplicity formula, we describe and enumerate the terms contributing a nonzero value to the multiplicity of a positive root $\mu$ in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$, which we denote $L(\tilde{\alpha})$, where $\tilde{\alpha}$ is the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. We prove that the number of terms contributing a nonzero value to the multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in $L(\tilde{\alpha})$ is given by the product $F_{i}\cdot F_{r-j+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. Using this result, we show that the $q$-multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in the representation $L(\tilde{\alpha})$ is precisely $q^{r-h(\mu)}$, where $h(\mu)=j-i+1$ is the height of the positive root $\mu$. Setting $q=1$ recovers the known result that the multiplicity of a positive root in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$ is one.
@@ 1 Introduction
For $r\geq 1$, the type $A_{r}$ Lie algebra, also known as the special linear algebra, is denoted $\mathfrak{sl}_{r+1}(\mathbb{C})=\{X\in M_{r+1}(\mathbb{C}):\text{Tr}(X)=0\}$. We recall that $\mathfrak{sl}_{r+1}(\mathbb{C})$ is a complex vector space with a bilinear product called the Lie bracket, defined by the commutator bracket $[X,Y]=XY-YX$ for $X,Y\in\mathfrak{sl}_{r+1}(\mathbb{C})$, which is skew-symmetric and satisfies the Jacobi identity. Given $e_{1},e_{2},\ldots,e_{r+1}$, the standard basis vectors of $\mathbb{R}^{r+1}$, for each $i\in[r]$, we define $\alpha_{i}=e_{i}-e_{i+1}$ We let $\Delta=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ be the set of simple roots, $\Phi^{+}=\Delta\cup\{\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}:1\leq i<j\leq r\}$ be the set of positive roots, and $\tilde{\alpha}=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{r}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$.
We recall that the Weyl group $W$ is the group generated by reflections, denoted as $s_{1},s_{2},\ldots,s_{r}$, through hyperplanes which are orthogonal to the simple roots $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$, respectively. The Weyl group of type $A_{r}$ is isomorphic to the symmetric group $\mathfrak{S}_{r+1}$. If $\sigma\in W$, then $\ell(\sigma)=k$ is the length of $\sigma$, which means that $k$ is the smallest nonnegative integer such that $\sigma$ can be expressed (not necessarily uniquely) as a product of $k$ reflections.
One area of interest in combinatorial representation theory is the computation involved in determining the multiplicity of a weight $\mu$ in a highest weight representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$ with highest weight $\lambda$. One way to compute this multiplicity is via Kostant’s weight multiplcity formula. For weights $\lambda$ and $\mu$, let the Kostant’s weight multiplicity formula to be
@@
where $\rho=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha$ and $\wp(\xi)$ is Kostant’s partition function, which counts the number of ways the weight $\xi$ can be written as a nonnegative integral sum of positive roots (Theorem on page 589 in [14]).
When calculating the Kostant’s weight multiplicity formula (1), every element in the Weyl group contributes a term to the multiplicity. We know the order of the Weyl group increases factorially as the rank of the Lie algebra increases, and it has been observed that many of the terms contribute trivially (i.e. give a zero term) to the computation [5]. To formulate a more efficient way to compute Kostant’s weight multiplicity formula, Harris gave the following definition.
Definition 1.1 (Definition 1.1 in [9]).
For $\lambda$ and $\mu$ integral weights of $\mathfrak{g}$, let the Weyl alternation set to be
@@
This set tells us the elements of the Weyl group for which $\sigma(\lambda+\rho)-\mu-\rho$ can be written as a nonnegative integral sum of positive roots. Knowing the Weyl alternation set allows one to reduce the weight multiplicity computation by only needing to consider those Weyl group elements that contribute nonzero terms to this computation.
An additional challenge in computing weight multiplicities is that there are no general formulas for Kostant’s partition function nor for its $q$-analog. We recall the following result of Lusztig, which defines the $q$-analog of Kostant’s weight multiplicity formula.
Definition 1.2 (Proposition 9.2 in [16]).
The $q$-analog of Kostant’s partition function is the polynomial-valued function
$\wp_{q}(\xi)=c_{0}+c_{1}q+c_{2}q^{2}+\cdots+c_{k}q^{k},$ (3)
@@
where $c_{i}$ equals the number of ways to write $\xi$ as a sum of exactly $i$ positive roots. Then the $q$-analog of Kostant’s weight multiplicity formula is given by
$m_{q}(\lambda,\mu)=\sum_{\sigma\in W}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\lambda+\rho)-\mu-\rho).$ (4)
Observe that evaluating equation (4) at $q=1$ recovers the multiplicity in equation (1).
@@
Work on characterizing and enumerating the Weyl alternation sets for certain pairs of weights, $(\lambda,\mu)$, has gathered some attention in the literature. This includes the following. If $\lambda$ and $\mu$ are integral weights of the Lie algebras $\mathfrak{sl}_{3}(\mathbb{C})$ and $\mathfrak{sl}_{4}(\mathbb{C})$, then the sets $\mathcal{A}(\lambda,\mu)$ are known in general [6, 15]. Harris considers $\tilde{\alpha}$, the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$ with $r\geq 1$, and $\mu=0$ and establishes the cardinality of the Weyl alternation set, $|\mathcal{A}(\tilde{\alpha},0)|$, is a Fibonacci number [9, Theorem 1.2]. Harris also states that if $\mu$ is a nonzero dominant weight, then the Weyl alternation set is either $\{1\}$ when $\mu=\tilde{\alpha}$ or empty otherwise. Chang, Harris, and Insko extend these results and show that when $\mu=0$ and $\lambda$ is the sum of all the simple roots of the simple Lie algebra $\mathfrak{so}_{2r+1}(\mathbb{C})$, the cardinality of the Weyl alternation set $\mathcal{A}(\lambda,0)$ is given by a Fibonacci number [4, Corollary 2.1 and 3.1]. While in the Lie algebras $\mathfrak{sp}_{2r}(\mathbb{C})$ and $\mathfrak{so}_{2r}(\mathbb{C})$, the cardinalities of the analogous Weyl alternation sets $\mathcal{A}(\lambda,0)$ are given by a multiple of a Lucas number [4, Corollary 4.1 and 5.1]. Harris, Insko, and Williams consider $\tilde{\alpha}$, the highest root of the classical Lie algebras, and $\mu=0$ and show that $|\mathcal{A}(\tilde{\alpha},0)|$ satisfies homogeneous linear recurrence relations [11]. In the type $B$ case, if $\mu=\varpi_{1}$ denotes the first fundamental root, then $|\mathcal{A}(\lambda,\varpi_{1})|$ is a Fibonacci number [11, Corollary 7.4]. They also consider the case where $\mu$ is a dominant positive root and give the multiplicity of these weights in the adjoint representation of the classical Lie algebras. Berenshteĭn and Zelevinskiĭ [1, Theorem 1.3] provide a comprehensive list of pairs of weights $(\lambda,\mu)$ of the classical Lie algebras and the exceptional Lie algebra $G_{2}$ for which $m(\lambda,\mu)=1$. Harris, Rahmoeller, Schneider, and Simpson consider certain such pairs, determine $\mathcal{A}(\lambda,\mu)$, and show that the $q$-multiplicity is a power of $q$ [12]. For all pairs of weights that Berenshteĭn and Zelevinskiĭ provide, Harris et. al conjecture that $m_{q}(\lambda,\mu)$ is a power of $q$ [12, Conjecture 7], and as of yet, this conjecture remains an open problem.
@@
As noted above, past work fixed $\lambda$ to be the highest root or the sum of the simple roots of a classical Lie algebra, while $\mu$ was either 00 or a dominant positive root. In this work, we focus our study on $\mathfrak{sl}_{r+1}(\mathbb{C})$, fix $r\geq 1$, let $\tilde{\alpha}$ be the highest root, and consider $\mu\in\Phi^{+}$. In this way, $\mu$ may be a nondominant weight. It is known that whenever $\mu$ is a root, $m(\tilde{\alpha},\mu)=1$. Hence, when $\mu$ is dominant, the pair $(\lambda,\mu)$ is given by Berenshteĭn and Zelevinskiĭ [1]. Moreover, we recall that the multiplicity is invariant under the action of the Weyl group. Namely, $m(\tilde{\alpha},\mu)=m(\tilde{\alpha},\sigma(\mu))$ for all $\sigma\in W$ [7, Proposition 3.1.20]. However, when $\mu\in\Phi^{+}$ is nondominant, a characterization and enumeration of the elements in $\mathcal{A}(\tilde{\alpha},\mu)$ was unknown. This paper contributes to this literature by proving the following results.
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then
where $F_{n}$ denotes the $n^{\text{th}}$ Fibonacci number [17, A000045].
Given the characterization and enumeration of the terms contributing nontrivially to the $q$-analog of Kostant’s weight multiplicity formula, we prove the following result using purely combinatorial techniques.
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then
@@
Observe that setting $q=1$ in Theorem 4.1 recovers the known result that the multiplicity of a positive root in the adjoint representation is equal to one [7, Theorem 2.4.1].
We remark that the work in [2, 3, 13], together imply that when $\tilde{\alpha}$ is the highest root, $\mu$ is a dominant weight, and $w$ is a Weyl group element, the jump polynomial associated to the weight $w(\mu)$ is equal to a power of $q$ times the $q$-analog of Kostant’s weight multiplicity formula $m_{q}(\lambda,\mu)$. In such a case, $m(\tilde{\alpha},\mu)=1$, and $m_{q}(\tilde{\alpha},\mu)$ is a power of $q$. Thus, providing an alternative proof of Theorem 4.1. However, the proofs presented in the current are purely combinatorial.
This work is organized as follows: In Section 2, we present the proof to Theorem 2.1 showing that the cardinality of the Weyl alternation set is a product of Fibonacci numbers. In Section 3, we establish results related to the $q$-analog of Kostant’s partition function. In Section 4, we present a proof to Theorem 4.1 showing that the $q$-analog of Kostant’s weight multiplicity formula of a positive root in the adjoint representation simplifies down to a power of $q$.
Acknowledgements. The author thanks her adviser Dr. Pamela E. Harris for the introduction to the problem and feedback on an earlier draft. She also thanks Dr. Jeb Willenbring for insightful comments throughout the progression of this manuscript. The author gratefully acknowledges funding from the University of Wisconsin - Milwaukee Department of Mathematical Sciences in support of this research.
@@ 2 Weyl Alternation Sets
In [4], they consider the case where $\lambda$ is the highest root and $\mu=0$. We still let $\lambda$ be the highest root and denote it as $\tilde{\alpha}$. However, we let $\mu\in\Phi^{+}$. The following result characterizes the elements in the set $\mathcal{A}(\tilde{\alpha},\mu)$. The following theorem is closely related to Theorem 2.1 in [4], and we adapt the proof techniques for our purpose.
@@
Fix $1\leq i\leq j\leq r$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Then $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
@@
($\Rightarrow$) We proceed by contradiction and assume $\sigma\neq 1$ or $\sigma\neq s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$. Then $\sigma$ must contain either $s_{1}$, $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
@@
since $-\alpha_{1}+\alpha_{2}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r}$ cannot be written as a sum of positive roots since the coefficient for $\alpha_{1}$ is negative. Thus, we get a contradiction, and $s_{1}\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. By [11, Proposition 3.4], if $\sigma\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$, then neither is any $\sigma^{\prime}\in W$ that contains $\sigma$ in its reduced word expression. Thus, $\sigma\in W$ containing $s_{1}$ in its reduced word expression cannot be in $\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. This reaches a contradiction to the assumption that $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$.
A similar argument is made for when $\sigma$ contains either $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
@@
Notice that $\{i,i+1,\ldots,j\}\cap\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}=\emptyset$. If $x\in\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (5), the coefficient of $\alpha_{x}$ is exactly 0. If $x\notin\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (5), the coefficient of $\alpha_{x}$ is exactly 1. Hence, $\tilde{\alpha}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})-\sum_{h=1}^{t}\alpha_{\ell_{h}}$, can be written as a sum of positive roots. Therefore, $s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. ∎
@@
We utilize the following enumerative result in later results.
@@
The number of ways to select a subset of $[n]:=\{1,2,3,\ldots,n\}$ consisting of nonconsecutive numbers is given $F_{n+2}$, where $F_{i}$ is the $i^{\text{th}}$ Fibonacci number.